OpenStudy (anonymous):

Consider the 639 N weight held by two cables shown below. The left-hand cable had ten- sion T2 and makes an angle of 53 with the ceiling. The right-hand cable had tension T1 and makes an angle of 49 with the ceiling.a) What is the tension in the cable labeled T1 slanted at an angle of 49? b.) And T2 slanted at 53...

OpenStudy (anonymous):

ok so I have actually worked it out.... once again i need a checker to make sure i did everything right....

OpenStudy (anonymous):

Need help?

OpenStudy (anonymous):

ok I got Im calling The straight Vector starting at the points where T1 and T2 meet T3. And t3 = m * g = 639N * 9.8 m/s^2= this = 6262.2 N* m /s^2 (im not for sure about the units thats what i need help on) Fx = -T1 cos 53 + T2cos 49 = 0 Fy = T1 sin 53 + T2 sin 49 - 6262.2 = 0

OpenStudy (anonymous):

o kk

OpenStudy (anonymous):

.8 T1 + .75T2 = 6262.2 T2 = .6T1/.66

OpenStudy (anonymous):

look at the link krystal see if that helps u kk?

OpenStudy (anonymous):

t1=4231.22 then plug that in to the other formula T2=3836.31

OpenStudy (anonymous):

Hey krystal. Your methodology looks solid. Remember that a Newton has units\[\rm [N] = \left [ kg \cdot m \over s^2 \right]\]Therefore, you don't need to multiply 639N by gravity because in this case Newton is a measure of weight not mass.

OpenStudy (anonymous):

so gravity was not needed in this case....

OpenStudy (anonymous):

well its always needed :)

OpenStudy (anonymous):

Nope. \[{Newtons \over gravity} = mass\]

OpenStudy (anonymous):

so dont multiply divide.....ok

OpenStudy (anonymous):

No no no. We want units of Newtons. Sorry for the confusion. Remember that Newton is the unit of force, which is what we are considering here.

OpenStudy (anonymous):

okokokko. :) thanks

OpenStudy (anonymous):

When we say that something weight xxx N, it has already been multiplied by gravity. In other cases, we are given mass and we multiply by gravity to get the weight. Here that has been done for you.